# Quantum Mixing and Benjamini–Schramm Convergence of Hyperbolic Surfaces

### v1

@@ 4. Spectral data
In this section we prove a bound for spectral quantity
@@
under certains assumptions on $a,b,t,\tau,$ and $T$, given that
@@
The spectral quantity appears in the proof of Theorem 1.5 and there it contains the spectral data of the problem.
We state the proposition.
@@
where $0<\delta<\tfrac{2}{9}\,m$ and where $T\,\delta=\pi/2$. Then
@@
The proof consists of relatively straightforward integration and careful application of the relationships of the variables and coefficients. The simplicity follows from a convenient choice of $h_{t}$ and $h_{t,\tau}$.
@@
from below to prove the bound. $G_{t,\tau}$ is illustrated in Figure 1.
Figure 1. Contour plot of $G_{\tau,T}(a,b)$. Two pronounced “ridges” are visible along the lines $b=a+\tau$ and $b=a-\tau$. The ridges get pronounced as $T$ increases. For a fixed $\delta$, $G_{\tau,t}(a,b)$ has a positive lower bound when $a-b-\tau\in(-\delta,\delta)$. Notice also how the ridges decay as $a$ and $b$ grow.
By using the definitions of $h_{t}$ and $h_{t,\tau}$, we write $G_{\tau,T}(a,b)$ as
@@
By using trigonometric sum-to-product identities, we have that $\cos(t\tau)\sin(ta)$ equals
$\frac{1}{2}\left(\sin\left(t(\tau+a)\right)-\sin\left(t(\tau-a)\right)\right),$
and further that $\cos(t\tau)\sin(ta)\sin(tb)$ equals
$\frac{1}{4}\left(\cos(t(a-b-\tau))+\cos(t(a-b+\tau)-\cos(t(a+b-\tau))-\cos(a+b+\tau))\right).$
Using this equality, we have that Expression (4.1) equals
$\frac{1}{4Tab}\left|\int_{0}^{T}\cos(t(a-b-\tau))+\cos(t(a-b+\tau)-\cos(t(a+b-\tau))-\cos(a+b+\tau))\,dt\right|.$
The sine terms of the integral can all be integrated separately and thus the integral yields the following equal expression:
$\frac{1}{4Tab}\left|\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}-\frac{\sin(T(a+b-\tau))}{a+b-\tau}-\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|$
Applying the reverse triangle inequality yields a following lower bound for the previous expression:
@@
We proceed to bound the absolute value terms separately. Consider the first absolute value term of the previous expression. Consider function
$x\mapsto\frac{\sin(T\,x)}{x}.$
It is positive when $x\in\left(-\pi/T,\pi/T\right)$, and it also decreases as $|x|$ grows in this interval. Because $T\delta=\pi/2$ and $a-b-\tau\in(-\delta,\delta)$, we can apply the observations regarding the previously studied function to get that
@@
We also notice that
$\frac{\sin(Tx)}{x}>0$
when $x\in\left(-\pi/T,\pi/T\right)$, which means that for every $x\in\mathbb{R}$ we have that
@@
The previous two observations mean that
$\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}$
is necessarily positive and hence
$\left|\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}\right|=\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}>\frac{1}{2\delta}.$
Next, consider the second and third absolute value terms of Expression (4.2). We observe that
@@
where we used the reverse triangle inequality together with the fact that $|a-b-\tau|$ is smaller than $\delta$ and that $2m$ is larger than $9\delta$. Using these observations gives that
$\left|\frac{\sin(T(a+b-\tau))}{a+b-\tau}\right|\leq\frac{1}{|a+b-\tau|}<\frac{1}{8\delta}.$
A similar approach yields that
$\left|\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|<\frac{1}{8\delta}.$
These two observations mean that
@@
With all the absolute value terms bounded, Expression (4.2) has the following lower bound:
$\frac{1}{4Tab}\,\frac{1}{4\delta}=\frac{1}{16\frac{\pi}{2}ab}=\frac{1}{8\pi ab}.$
In the previous we used that $T\delta=\pi/2$.
Combining these bounds gives that
$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|\ \geq\ \frac{1}{4Tab}\cdot\frac{1}{4\delta}\ =\ \frac{1}{16\,\frac{\pi}{2}\,ab}\ =\ \frac{1}{8\pi\,ab}.$
Therefore,
$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|^{-1}=G_{\tau,T}^{-1}<8\pi ab,$
as claimed. ∎
@@ 5. Propagators
the geometric data of the problem that appears in the proof of Theorem 1.5, it is necessary to analyze the kernels of the hyperbolic wave propagator $P_{t}$ and a slightly mofified version of it, $P_{t,\tau}$, for these kernels are then used to open the Hilbert-Schmidt norms. Although the hyperbolic wave propagator is a natural choice, it exhibits greater thechnical complexity than, for instance, the ball averaging operator (cf. [LS17]). As mentioned earlier, this choice makes the specral side easier.
We state a proposition giving a convenient form for the integral kernels of $P_{t}$ and $P_{t,\tau}$.
@@
be defined on a compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$ with $\Delta_{X}$ being the Laplace-Beltrami operator and where
@@
Then $P_{t}$ admits integral kernel
@@
and $P_{t,\tau}$ admits integral kernel $\cos(t\tau)\,K_{t}^{\Gamma}$.
Given that
$P_{t}=h_{t}\left(\sqrt{-\Delta_{X}-\frac{1}{4}}\right)$
and that $\tilde{P}_{t}$ is the operator with the integral kernel $K_{t}^{\Gamma}$ which was presented in the statement of the proposition, the proposition can be proved by showing that inner product $\langle P_{t}f,g\rangle$ equals inner product $\langle\tilde{P}_{t}f,g\rangle$ where $f,g\in C^{\infty}(X)$ as $P_{t}$ and $\tilde{P}_{t}$ are bounded operators on $L^{2}(X)$ with $X$ being compact. The core idea is representing $P_{t}$ as a limit of propagators $P_{t}^{(\varepsilon)}$ that have a convenient explicit formulae for their integral kernels.
The proof relies on multiple auxiliary results presented after the proof of Proposition 5.1 for clarity’s sake
@@
The assertion for $P_{t,\tau}$ follows directly from the assertion for $P_{t}$ since $P_{t,\tau}=\cos(t\tau)P_{t}$. Hence it suffices to consider $P_{t}$ and $K_{t}$.
Due to way $P_{t}$ is defined, it is bounded and has property
@@
where $\psi_{j}$ is an eigenfunction of the Laplacian on $X$ with eigenvalue $\lambda_{j}$ such that $\rho_{j}=\sqrt{\lambda_{j}-1/4}$. Define regularized operators for $\varepsilon>0$:
$P_{t}^{(\varepsilon)}=h_{t}^{(\varepsilon)}\Big(\sqrt{-\Delta-\tfrac{1}{4}}\Big),\quad h_{t}^{(\varepsilon)}(x)=\frac{\sin(tx)}{x}\,e^{-\frac{\varepsilon^{2}x^{2}}{2}}.$
Due to way $P_{t}^{(\epsilon)}$ is defined, it is bounded and has property
@@
Since $h_{t}^{(\varepsilon)}$ converges to $h_{t}$ pointwise as $\varepsilon$ shrinks, we have that $P_{t}^{(\varepsilon)}f$ converges to $P_{t}f$ pointwise as $\varepsilon$ shrinks for every $f\in C^{\infty}(X)$.
Proposition 5.2 ensures the existence of integral operator $Q_{t}^{(\varepsilon)}$ on $X$ with integral kernel
@@
Lemma 5.3 allows us to rewrite $k_{t}^{(\varepsilon)}$ in a more tractable form so that
where
@@
denotes the standard Gaussian mollifier. Since an operator on $X$ is uniquely determined by its action on the orthonormal basis, it follows that $Q_{t}^{(\varepsilon)}=P_{t}^{(\varepsilon)}$.
Let $\tilde{P}_{t}$ denote an integral operator on $X$ with kernel $K_{t}^{\Gamma}$. To show that $\tilde{P}_{t}$ is well-defined and bounded, we use the Schur test (see [HS12]), so it suffices to verify that for all $x\in X$
$\int_{X}K_{t}(x,y)\,dy$
is finite as $K_{t}^{\Gamma}(x,y)$ is symmetric in $x$ and $y$. Open the previous expression using the definition of $K_{t}$:
$\frac{1}{2\sqrt{2}\pi}\,\int_{X}\sum_{\gamma\in\Gamma}\frac{\bm{1}_{t>d(x,\gamma y)}}{\sqrt{\cosh(t)-\cosh(d(x,\gamma y))}}dy.$
Uniting the integral and the sum into an integral over the universal cover gives the following equal expression:
$\frac{1}{2\sqrt{2}\pi}\int_{\mathbb{H}}\frac{\bm{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}dy.$
Writing this using polar coordinates around $x$ yields
@@
where coefficient $2\pi$ appeared first from the polar coordinates since there was no dependency on the direction, and the equality followed from computing the integral. The result is finite and $\tilde{P}$ is well-defined and bounded.
To show that $P_{t}$ has integral kernel $K_{t}^{\Gamma}$, as claimed in the statement of the proposition, it suffices to verify that for every $f,g\in C^{\infty}(X)$
$\langle P_{t}f,g\rangle-\langle\tilde{P}_{t}f,g\rangle=0\iff\left|\langle P_{t}f,g\rangle-\langle\tilde{P}_{t}f,g\rangle\right|=0,$
since $P_{t}$ and $\tilde{P}_{t}$ are bounded operators on $L^{2}(X)$ where $X$ is compact. To show this, start by considering the absolute value term of the previous expression. It can be written as
$\left|\int_{X}g(x)^{*}\big(P_{t}f(x)-\tilde{P}_{t}f(x)\big)dx\right|,$
where we opened the inner product and united the two integrals. By pulling the absolute values inside the integral and then bounding $|g|$ with the $\sup$ norm, we get the following upper bound for the previous expression:
@@
Focus on the absolute value integrand appearing in the previous expression. We can write it in the following form by writing $P_{t}f$ using the limit and then by writing the operator action using the kernels:
@@
Writing $K_{t}^{\Gamma}$ and $K_{t}^{(\varepsilon)}$ using their sum representations over $\Gamma$ and then uniting the integral and the sum into an integral over the universal cover yields the following equal expression, where we also view $f$ as $\Gamma$ periodic function:
@@
By pulling the absolute values inside the integral and bouding $|f|$ with the $\sup$ norm, we get the following upper bound:
$\|f\|_{\infty}\lim_{\varepsilon\to 0}\int_{\mathbb{H}}\left|k_{t}^{(\varepsilon)}(x,y)f(y)-K_{t}(x,y)\right|dy.$
Focus on the integral in the previous expression. By opening $k_{t}$ and $K_{t}$, the integral can be written as
$\frac{1}{2\sqrt{2}\pi}\int_{\mathbb{H}}\left|\left(\int_{\mathbb{R}}\frac{\left(\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(-t-v)\right)\,\bm{1}_{v>d(x,y)}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}dv\right)-\frac{\bm{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}\right|dy.$
By defining
$\chi_{d(x,y)}(t)=\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}},$
we may write the preceding expression as
@@
where $\ast_{\mathbb{R}}$ denotes the convolution on $\mathbb{R}$. The additional term $\chi_{d(x,y)}(-t)$ equals 0 so adding it had no effect on the equality. Using the triangle inequality, the previous expression is bounded by
$\frac{1}{2\sqrt{\pi}}\lim_{\varepsilon\to 0}\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\chi_{d(x,y)}(t)\right|+\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(-t)-\chi_{d(x,y)}(-t)\right|\,dy.$
Finally, Lemma 5.4 guarantees that $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ converges to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$. The line of reasoning leading to this cocnlusion means that
$\langle P_{t}f,g\rangle=\langle\tilde{P}_{t}f,g\rangle,$
@@
The remaining section contains the lemmas and propositions used in the proof of Proposition 5.1. First, we record a slightly modified formulation of Proposition 5 in [Mar04]. It gives a family of operators with explicit formulae for their integral kernels on the compact connected hyperbolic surface. Especially, these operators share the eigenfunctions with the Laplacian of the surface and of which eigenvalues have a representation using the eigenvalues of the Laplacian.
@@
Let $h:\{x+iy:|y|<C\}\to\mathbb{C}$ satisfy:
  1. i.
@@
  3. iii.
    $h(x+iy)$ decays superpolynomially as $|x|\to\infty$,
@@
where $\psi_{j}$ is a Laplacian eigenfunction on $X$ with eigenvalue $\lambda_{j}$ and $\rho_{j}=\sqrt{\lambda_{j}-1/4}$.
Next, we derive a convenient expression for $k_{t}^{(\varepsilon)}$, an integral kernel that is crucial in establishing a convenient formula for the integral operator $P_{t}$.
@@
The proof is based on noting that there is a Fourier transform inside the expression used to define $k_{t}^{(\varepsilon)}$ in the statement. Then by using the inverse Fourier transform of $\phi_{\varepsilon}$ and the Fourier transform of $s_{t}(x)=x^{-1}\,\sin(t\,x)$ (in the distributional sense) together with the Fourier convolution identity, we acquire the wanted conclusion.
@@
Applying Fubini’s theorem to interchange the order of integration, we obtain that $k_{t}^{(\varepsilon)}$ equals
@@
and hence the Inverse Fourier transform is
$F^{-1}f(x)=\frac{1}{2\pi}\int_{\mathbb{R}}f(y)\,e^{ixy}dy.$
@@
By using the Fourier convolution identity and the Fourier transform of sine we observe that the inner integral of Expression (5.1) equals
@@
where $s_{t}$ denotes function $x\mapsto\sin(tx)$. Substituting this into Expression (5.1) yields the following equal expression:
$\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(v-t)-\phi_{\varepsilon}(v+t)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv=\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(t+v)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv,$
where the equality followed using the symmetry of the Gaussian mollifier.
Finally, the following lemma establishes convergence of $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$. Interestingly, the convolution here is over the real line but the convergence is studied in the $\mathbb{H}$.
@@
Then map $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ converges to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$ for every $t\in\mathbb{R}$ as $\varepsilon\to 0$, where $\ast_{\mathbb{R}}$ denotes the convolution on $\mathbb{R}$.
In the proof we show that
@@
The idea is to present this integral as a convolution $\phi_{\varepsilon}\ast_{\mathbb{R}}A(t)$, where $A$ is a continuous function with $A(t)=0$. Then the conclusion can be drawn using the properties of the convolution.

### v2

@@ 4. Spectral data
In this section, we prove a bound for the spectral quantity
@@
under certain assumptions on $a,b,t,\tau,$ and $T$, given that
@@
This spectral quantity appears in the proof of Theorem 1.5, where it encodes the spectral data of the main problem. We state the proposition.
@@
where $\delta\in(0,\tfrac{2}{9}m)$. Then for $T\delta=\tfrac{\pi}{2}$ we have that
@@
The proof consists of straightforward integration and careful application of the relationships among the variables. The simplicity follows from the convenient choices of $h_{t}$ and $h_{t,\tau}$.
@@
from below. $G_{\tau,T}$ is illustrated in Figure 1.
Figure 1. Contour plot of $G_{\tau,T}(a,b)$. Two pronounced “ridges” are visible along the lines $b=a+\tau$ and $b=a-\tau$. The ridges become more pronounced as $T$ increases and decay as $a$ and $b$ increase.
Using the definitions of $h_{t}$ and $h_{t,\tau}$, we expand $G_{\tau,T}(a,b)$ to
@@
Applying trigonometric sum-to-product identities, we have
$\cos(t\tau)\sin(ta)=\frac{\sin\left(t(\tau+a)\right)-\sin\left(t(\tau-a)\right)}{2},$
and hence
$\cos(t\tau)\sin(ta)\sin(tb)=\frac{\cos(t(a-b-\tau))+\cos(t(a-b+\tau))-\cos(t(a+b-\tau))-\cos(t(a+b+\tau))}{4}.$
Therefore, Expression (4.1) equals
$\frac{1}{4Tab}\left|\int_{0}^{T}\cos(t(a-b-\tau))+\cos(t(a-b+\tau))-\cos(t(a+b-\tau))-\cos(t(a+b+\tau))\,dt\right|.$
Integrating each cosine term separately gives
$\frac{1}{4Tab}\left|\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}-\frac{\sin(T(a+b-\tau))}{a+b-\tau}-\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|.$
Applying the reverse triangle inequality yields a lower bound:
@@
We now bound the absolute value terms separately. Consider the first absolute value. The function
$x\mapsto\frac{\sin(T\,x)}{x}$
is positive when $x\in\left(-\frac{\pi}{T},\frac{\pi}{T}\right)$, and decreases as $|x|$ grows in this interval. Since $T\delta=\frac{\pi}{2}$ and $a-b-\tau\in(-\delta,\delta)$, we have
@@
Moreover, for all $x\in\mathbb{R}$:
@@
Thus,
$\frac{\sin(T(a-b-\tau))}{a-b-\tau}+\frac{\sin(T(a-b+\tau))}{a-b+\tau}>\frac{1}{2\delta}.$
Next, consider the second and third absolute value terms in Expression (4.2). Since
@@
$|a+b+\tau|\geq 2a-|a-b-\tau|>2m-\delta>8\delta.$
Hence,
$\left|\frac{\sin(T(a+b-\tau))}{a+b-\tau}\right|<\frac{1}{8\delta},\qquad\left|\frac{\sin(T(a+b+\tau))}{a+b+\tau}\right|<\frac{1}{8\delta},$
so that
@@
Combining these bounds, Expression (4.2) has the lower bound
$\frac{1}{4Tab}\,\frac{1}{4\delta}=\frac{1}{16\frac{\pi}{2}ab}=\frac{1}{8\pi ab},$
where we used $T\delta=\frac{\pi}{2}$. Hence,
$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|>\frac{1}{8\pi ab},$
and therefore
$\left|\frac{1}{T}\int_{0}^{T}h_{t,\tau}(a)\,h_{t}(b)\,dt\right|^{-1}=G_{\tau,T}^{-1}<8\pi ab.$
@@ 5. Propagators
the geometric data of the problem that appears in the proof of Theorem 1.5, it is necessary to analyze the kernels of the hyperbolic wave propagator $P_{t}$ and a slightly modified version of it, $P_{t,\tau}$, since these kernels are then used to bound the Hilbert–Schmidt norms. Although the hyperbolic wave propagator is a natural choice, it exhibits greater technical complexity than, for instance, the ball averaging operator (cf. [LS17]); the choice is deliberate as it makes the spectral side easier. We state a proposition giving a convenient form for the integral kernels of $P_{t}$ and $P_{t,\tau}$.
@@
be defined on a compact connected hyperbolic surface $X=\Gamma\setminus\mathbb{H}$, with $\Delta_{X}$ being the Laplace–Beltrami operator, and where
@@
Then $P_{t}$ admits the integral kernel $K_{t}^{\Gamma}$ defined as
@@
and $P_{t,\tau}$ admits the integral kernel $\cos(t\tau)\,K_{t}^{\Gamma}$.
Given that $\tilde{P}_{t}$ is the integral operator with integral kernel $K_{t}^{\Gamma}$, the proposition can be proved by showing that $\langle P_{t}f,g\rangle=\langle\tilde{P}_{t}f,g\rangle$ for $f,g\in C^{\infty}(X)$, since $P_{t}$ and $\tilde{P}_{t}$ are bounded operators on $L^{2}(X)$. The core idea of showing the equality of the inner products is representing $P_{t}$ as a limit of propagators $P_{t}^{(\varepsilon)}$ that have convenient explicit formulae for their integral kernels. The proof relies on multiple auxiliary results presented after the proof of Proposition 5.1.
@@
The assertion for $P_{t,\tau}$ follows directly from the assertion for $P_{t}$ since $P_{t,\tau}=\cos(t\tau)P_{t}$. Hence, it suffices to consider $P_{t}$ and $K_{t}^{\Gamma}$.
Due to the way $P_{t}$ is defined, it is bounded and has the property
@@
where $\psi_{j}$ is an eigenfunction of $\Delta_{X}$ with eigenvalue $\lambda_{j}$ such that $\rho_{j}=\sqrt{\lambda_{j}-\frac{1}{4}}$. Define regularized operators for $\varepsilon>0$:
$P_{t}^{(\varepsilon)}=h_{t}^{(\varepsilon)}\Big(\sqrt{-\Delta_{X}-\tfrac{1}{4}}\Big),\quad h_{t}^{(\varepsilon)}(x)=\frac{\sin(tx)}{x}\,e^{-\frac{\varepsilon^{2}x^{2}}{2}}.$
Due to the way $P_{t}^{(\varepsilon)}$ is defined, it is bounded and has the following property:
@@
Since $h_{t}^{(\varepsilon)}$ converges to $h_{t}$ pointwise as $\varepsilon\to 0$, we have that $P_{t}^{(\varepsilon)}f$ converges to $P_{t}f$ pointwise for every $f\in C^{\infty}(X)$ due to spectral theory.
Proposition 5.2 ensures the existence of an integral operator $Q_{t}^{(\varepsilon)}$ on $X$ with integral kernel $K_{t}^{(\varepsilon)}$ defined as
@@
Lemma 5.3 allows us to rewrite $k_{t}^{(\varepsilon)}$ in a more tractable form:
$k_{t}^{(\varepsilon)}(x,y)=\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(-t-v)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv.$
Here
@@
is the standard Gaussian mollifier. Since a bounded operator is uniquely determined by its action on an orthonormal basis, it follows that $Q_{t}^{(\varepsilon)}=P_{t}^{(\varepsilon)}$.
Let $\tilde{P}_{t}$ denote the integral operator on $X$ with kernel $K_{t}^{\Gamma}$. To show that $\tilde{P}_{t}$ is well-defined and bounded, we use the Schur test (see [HS78]). It suffices to verify that for all $x\in X$,
$\int_{X}K_{t}^{\Gamma}(x,y)\,dy<\infty,$
since $K_{t}^{\Gamma}(x,y)$ is symmetric in $x$ and $y$. We open the integral using the definition of $K_{t}^{\Gamma}$:
$\frac{1}{2\sqrt{2}\pi}\,\int_{X}\sum_{\gamma\in\Gamma}\frac{\mathbf{1}_{t>d(x,\gamma y)}}{\sqrt{\cosh(t)-\cosh(d(x,\gamma y))}}dy.$
Uniting the integral and the sum into an integral over the universal cover gives:
$\frac{1}{2\sqrt{2}\pi}\int_{\mathbb{H}}\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}dy.$
Writing this using polar coordinates around $x$ yields:
@@
where the coefficient $2\pi$ arises from the polar coordinates, and the equality follows from computing the integral. The result is finite, and $\tilde{P}_{t}$ is well-defined and bounded.
To show that $P_{t}$ has integral kernel $K_{t}^{\Gamma}$, as claimed in the proposition, it suffices to verify that for every $f,g\in C^{\infty}(X)$,
$\left|\langle P_{t}f,g\rangle-\langle\tilde{P}_{t}f,g\rangle\right|=0,$
since $P_{t}$ and $\tilde{P}_{t}$ are bounded operators on $L^{2}(X)$ where $X$ is compact. To show this, consider the absolute value:
$\left|\int_{X}g(x)^{*}\big(P_{t}f(x)-\tilde{P}_{t}f(x)\big)dx\right|.$
By moving the absolute value inside the integral and bounding $|g|$ by its supremum norm, we obtain the upper bound:
@@
We focus on the absolute value inside the previous integral. Writing $P_{t}f$ using the limit and expressing the operator action via kernels, we obtain
@@
Expressing the kernels using the sum over $\Gamma$ and combining the sum with the integral over $X$ into an integral over the hyperbolic plane yields the following equivalent expression:
@@
Here, $f$ is viewed as a $\Gamma$-periodic function on $\mathbb{H}$. By moving the absolute value inside and bounding $|f|$ by its supremum norm, we obtain the upper bound:
$\|f\|_{\infty}\lim_{\varepsilon\to 0}\int_{\mathbb{H}}\left|k_{t}^{(\varepsilon)}(x,y)-K_{t}(x,y)\right|dy.$
Focus on the limit of the integral. By opening $k_{t}$ and $K_{t}$, the limit can be written as
$\lim_{\varepsilon\to 0}\frac{1}{2\sqrt{2}\pi}\int_{\mathbb{H}}\left|\left(\int_{\mathbb{R}}\frac{(\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(-t-v))\mathbf{1}_{v>d(x,y)}}{\sqrt{\cosh(v)-\cosh(d(x,y))}}dv\right)-\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}\right|dy.$
Define
$\chi_{d(x,y)}(t)=\frac{\mathbf{1}_{t>d(x,y)}}{\sqrt{\cosh(t)-\cosh(d(x,y))}}.$
Using this, the previous integral expression becomes
@@
where $\ast_{\mathbb{R}}$ denotes convolution on $\mathbb{R}$. The term $\chi_{d(x,y)}(-t)$ equals 0, so its addition has no effect. Using the triangle inequality, the expression is bounded by
$\frac{1}{2\sqrt{2}\pi}\lim_{\varepsilon\to 0}\int_{\mathbb{H}}\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)-\chi_{d(x,y)}(t)\right|+\left|\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(-t)-\chi_{d(x,y)}(-t)\right|dy.$
Finally, Lemma 5.4 guarantees that $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ converges to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$. Hence, the limit equals 0, and
$\left|\langle P_{t}f,g\rangle-\langle\tilde{P}_{t}f,g\rangle\right|=0,$
@@
The remaining part of this section contains lemmas and propositions used in the proof of Proposition 5.1. First, we record a slightly modified formulation of [MAR12, Prop. 5]. It gives a family of operators with explicit formulae for their integral kernels on a compact connected hyperbolic surface. These operators share the eigenfunctions with the Laplace–Beltrami operator of the surface, and their eigenvalues have a representation in terms of the Laplace–Beltrami eigenvalues.
@@
Let $h:\{x+iy\in\mathbb{C}\ :\ |y|<C\}\to\mathbb{C}$ satisfy the following properties, where $C$ is a positive constant:
@@
  • $h(x+iy)$ decays superpolynomially as $|x|\to\infty$.
@@
where $\psi_{j}$ is a Laplacian eigenfunction on $X$ with eigenvalue $\lambda_{j}$ and $\rho_{j}=\sqrt{\lambda_{j}-\tfrac{1}{4}}$.
In the next lemma, we derive a convenient expression for $k_{t}^{(\varepsilon)}$, which is an integral kernel crucial in establishing a tractable formula for the kernel of the integral operator $P_{t}$. More specifically, $k_{t}^{(\varepsilon)}$ is used to approximate the integral kernel of $P_{t}$. We state the lemma.
@@
The proof is based on noting that there is a Fourier transform inside the expression used to define $k_{t}^{(\varepsilon)}$. By using the inverse Fourier transform of $\phi_{\varepsilon}$ and the Fourier transform of $s_{t}(x)=x^{-1}\,\sin(tx)$ (in the distributional sense) together with the Fourier convolution identity, we obtain the desired conclusion.
@@
By applying Fubini’s theorem to interchange the order of integration, we obtain that $k_{t}^{(\varepsilon)}$ equals
@@
and the inverse Fourier transform is
$\mathcal{F}^{-1}f(x)=\frac{1}{2\pi}\int_{\mathbb{R}}f(y)\,e^{ixy}dy.$
@@
The inner integral in Expression (5.1) equals
@@
where the first equality follows from the convolution identity, and the second equality uses
$\mathcal{F}(s_{t})(x)=\pi i\bigl(\delta(x+t)-\delta(x-t)\bigr),$
with $s_{t}(x)=\sin(tx)$ and $\delta$ denoting the Dirac distribution.
Substituting this into Expression (5.1) gives:
$k_{t}^{(\varepsilon)}(x,y)=\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(v-t)-\phi_{\varepsilon}(v+t)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv=\frac{1}{2\sqrt{2}\pi}\int_{d(x,y)}^{\infty}\frac{\phi_{\varepsilon}(t-v)-\phi_{\varepsilon}(t+v)}{\sqrt{\cosh(v)-\cosh(d(x,y))}}\,dv,$
where the equality follows from the symmetry of the Gaussian mollifier. ∎
Finally, the following lemma establishes the convergence of $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$. Interestingly, the convolution is over the real line, but the convergence is studied in the hyperbolic plane.
@@
Then the map $y\mapsto\phi_{\varepsilon}\ast_{\mathbb{R}}\chi_{d(x,y)}(t)$ converges to $y\mapsto\chi_{d(x,y)}(t)$ in $L^{1}(\mathbb{H})$ for every $t\in\mathbb{R}$ as $\varepsilon\to 0$, where $\ast_{\mathbb{R}}$ denotes convolution on $\mathbb{R}$.
In the proof, we show that
@@
The idea is to present this integral as a convolution $\phi_{\varepsilon}\ast_{\mathbb{R}}A_{t}$, where $A_{t}$ is a continuous function with $A_{t}(t)=0$. Then the conclusion follows from the properties of convolution.